Three brothers, 3 years apart totaling 33 years. What's the age of the oldest brother ?
Answer by lenny460(1073)

Three brothers, 3 years apart totaling 33 years. What's the age of the oldest brother ?
Let their ages be n + (n + 3) + (n + 6) = 33
Therefore we now have:
3n + 9 = 33
3n + 9 - 9 = 33 - 9
3n = 24
n = 8
n + 3 = 11 and
n + 6 = 14

Answer: The oldest is 14 years old.



Lennox Obuong
Algebra Tutor